### v1

@@ Abstract
The argument sufficiency assessment task aims to determine if the premises of a given argument support its conclusion. To tackle this task, existing works often train a classifier on data annotated by humans. However, annotating data is laborious, and annotations are often inconsistent due to subjective criteria. Motivated by the probability of sufficiency (PS) definition in the causal literature, we propose Casa, a zero-shot causality-driven argument sufficiency assessment framework. PS measures how likely introducing the premise event would lead to the conclusion, when both the premise and conclusion events are absent. To estimate this probability, we propose to use large language models (LLMs) to generate contexts that are inconsistent with the premise and conclusion, and revise them by injecting the premise event. Experiments on two logical fallacy detection datasets demonstrate that Casa accurately identifies insufficient arguments. We further deploy Casa in a writing assistance application, and find that suggestions generated by Casa enhance the sufficiency of student-written arguments. Code and data are available at https://github.com/xxxiaol/CASA.
@@
[…]
@@ 3.1 Datasets
Correct Answer: Insufficient
Zero-shot Prompting (Tulu): Sufficient
@@ A.4 Additional Results
Correct Answer: Insufficient
Zero-shot Prompting (Llama2): Sufficient
@@ A.5 Details of the Writing Assistance Application
System
You are a helpful and educated assistant.
User
In each case, we will give you an argument and a model generated objection situation.

### v2

@@ Abstract
The argument sufficiency assessment task aims to determine if the premises of a given argument support its conclusion. To tackle this task, existing works often train a classifier on data annotated by humans. However, annotating data is laborious, and annotations are often inconsistent due to subjective criteria. Motivated by the definition of probability of sufficiency (PS) in the causal literature, we propose Casa, a zero-shot causality-driven argument sufficiency assessment framework. PS measures how likely introducing the premise event would lead to the conclusion when both the premise and conclusion events are absent. To estimate this probability, we propose to use large language models (LLMs) to generate contexts that are inconsistent with the premise and conclusion and revise them by injecting the premise event. Experiments on two logical fallacy detection datasets demonstrate that Casa accurately identifies insufficient arguments. We further deploy Casa in a writing assistance application, and find that suggestions generated by Casa enhance the sufficiency of student-written arguments. Code and data are available at https://github.com/xxxiaol/CASA.
@@
Xiao Liu1thanks:   Work done during visiting UCLA.,  Yansong Feng1and Kai-Wei Chang2 1Wangxuan Institute of Computer Technology, Peking University 2Computer Science Department, University of California, Los Angeles {lxlisa,fengyansong}@pku.edu.cn kwchang@cs.ucla.edu
@@ 3.1 Datasets
Correct Answer: Insufficient Zero-shot Prompting (Tulu): Sufficient
@@ 7 Conclusion
Acknowledgments
We thank Xueqing Wu, Fan Yin, Ashima Suvarna, Da Yin, and other UCLA-NLP lab members for their constructive comments. We thank the anonymous reviewers for their helpful discussions and suggestions. This research was supported in part by NSF #2331966.
@@ A.4 Additional Results
Correct Answer: Insufficient Zero-shot Prompting (Llama2): Sufficient
@@ A.5 Details of the Writing Assistance Application
System You are a helpful and educated assistant.
User In each case, we will give you an argument and a model generated objection situation.